OpenStudy (anonymous):

The population P in 1997 for a state is given along with r, it's annual percentage of continuous growth. P=33 millions, r=1.7% (a) Write the formula f(x) =P erx, where r is in decimal notation, that model the population in millions x years after 1997 (b)Estimate the population in 2015

OpenStudy (amistre64):

r = .017

OpenStudy (amistre64):

33,000,000 e^(.017*1000000(x))

OpenStudy (anonymous):

........

OpenStudy (anonymous):

33? Wth?

OpenStudy (anonymous):

Newton?

OpenStudy (anonymous):

33 million is written out there.

OpenStudy (anonymous):

that doesnt make sense

OpenStudy (anonymous):

The formula says Perx which is usually just an abbreviation for f(x)=Pe^(rx)

OpenStudy (amistre64):

its your question, not mine ;) im just the messenger....of doom!!

OpenStudy (amistre64):

do we have to use millions of years to find 2015?

OpenStudy (anonymous):

Dude, I have no clue, lol. That doesn;t make sense.

OpenStudy (amistre64):

why dont it make sense?

OpenStudy (amistre64):

large numbers are allowed to exist ;)

OpenStudy (amistre64):

or maybe it means: Write the formula f(x) =P erx, where r is in decimal notation, that model the population (in millions); x years after 1997

OpenStudy (amistre64):

33,000,000 e^(.017x)

OpenStudy (amistre64):

33 e^(.017x) seems more reasonable then.... if I parse the question correctly

OpenStudy (amistre64):

x = 18...(2015) 33 * (e^(.017 * 18)) = 44.8134161 million

OpenStudy (anonymous):

it wasn;t right....... :(

OpenStudy (amistre64):

which one, the 33 e^(.017x) ? cause that seems to be the most logical one

OpenStudy (anonymous):

correct answer was...........

OpenStudy (anonymous):

33e^0.017x

OpenStudy (anonymous):

33 million wasn't correct either ARGHH!

OpenStudy (amistre64):

:) whats the correct answer then

OpenStudy (anonymous):

44.8

OpenStudy (anonymous):

Ready for same problem, new numbers?

OpenStudy (amistre64):

I got that for b already

OpenStudy (anonymous):

you said 33 mil

OpenStudy (amistre64):

no, I said 33 million was how you set the problem up; not the answer; you gave a 2 parter. :)

OpenStudy (anonymous):

:(

OpenStudy (amistre64):

part 1: set up an equation to use with this info: 33 e^(.017x) part 2) use it to find the population in 18 years

OpenStudy (amistre64):

33 million was part of the setup; in part 1.

OpenStudy (anonymous):

The population P in 1991 for a state is given along with r, it's annual percentage of continuous growth. P=34 millions, r=2.3% (a) Write the formula f(x) =P erx, where r is in decimal notation, that model the population in millions x years after 1991 (b)Estimate the population in 2016

OpenStudy (amistre64):

PART A :) 34 e^(.023x) PART B ;) 2016-1991 = 25 34 * (e^(.023 * 25)) = 60.4224379

OpenStudy (anonymous):

that wasn't it! nooooooooooo

OpenStudy (anonymous):

I'm shocked!

OpenStudy (amistre64):

that should be it unless google lied to me in the math..

OpenStudy (anonymous):

That formula should be correct?

OpenStudy (amistre64):

60.4224379.... is the same answer I get on the calculator

OpenStudy (amistre64):

I used the formula they gave you and the numbers.... there isnt really a way to mess that up :) even for me...

OpenStudy (amistre64):

60.4 maybe is what they want

OpenStudy (anonymous):

I was trying the formula first the .......34 e^(.023x)

OpenStudy (anonymous):

for A

OpenStudy (amistre64):

i dont know how they want it formated; thats a computer program issue, but the formula itself is good

OpenStudy (anonymous):

aw nuts.

OpenStudy (anonymous):

it was 0.023

OpenStudy (amistre64):

....those bastards!.... they killed kenny

OpenStudy (anonymous):

second part was right! ;)